OpenStudy (anonymous):

2/y-15 = 3/5 Solve for y Please teach me how to do this?

OpenStudy (ranga):

Isolate the y. First add 15 to both sides and simplify the right Then turn both sides upside down (that is, take the reciprocal) multiply both sides by 2 to solve for y.

OpenStudy (anonymous):

Okay! Thank you! I wasn't sure if I could just times by the 2 first.

OpenStudy (ranga):

you are welcome. What did you get for y?

OpenStudy (anonymous):

Hold on a sec..almost got it

OpenStudy (anonymous):

31.2=y?

OpenStudy (anonymous):

nevermind 25/3 my bad

OpenStudy (ranga):

Only y is in the denominator right? Not (y-15) right?

OpenStudy (anonymous):

It's (y-15)

OpenStudy (ranga):

Oh then you should use parenthesis. Is this the equation?\[\Large \frac{ 2 }{ y-15} = \frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

Yes. Sorry haha

OpenStudy (ranga):

Then simply cross multiply: 2 * 5 = 3(y-15) divide by 3 10/3 = y - 15 add 15 10/3 + 15 = y 10/3 + 45/3 = y y = 55/3

OpenStudy (anonymous):

Oh okay

OpenStudy (ranga):

alright. When writing math equations horizontally without an editor always make use of parenthesis to let others know whether the whole thing is in the denominator, exponent, etc.

OpenStudy (anonymous):

Thank you so much :). I gotta go now though; I need to finish this homework soon :). Bye!

OpenStudy (ranga):

bye.